A traveling wave in a string is given by y=0.03 sin(2.2x-3.5t) where y and x are in metres and t is in seconds. Find the amplitude, wavelength, the frequency, the period and the of the wave?
1
Expert's answer
2020-11-10T06:59:39-0500

The traveling wave is described with the following equation:


"y = A\sin(kx-\omega t)"

where "A" is the amplitude, "\omega" is the angular frequency, and "k = 2\pi/\lambda" is the wave number, where "\lambda" is the wavelength. Comparing this general form with the given expression


"y = 0.03\sin(2.2x-3.5t)"

obtain:


"A = 0.03m\\
\lambda = 2\pi/k = 2\pi/2.2 \approx 2.86m\\
f = \dfrac{\omega}{2\pi} = \dfrac{3.5}{2\pi} \approx 0.57Hz\\
T = 1/f \approx 1.8s"

Answer. Amplitude 0.03m, wavelength 2.86m, frequency 0.57Hz, period 1.8s.